OpenStudy (anonymous):

Please help!!! solve for A,B,and C 2A+3B-4C=7 A -5B+2C=2 -7A-3B+C=-20

OpenStudy (anonymous):

solve it simultaneously by elimination method or substitution! :-)

OpenStudy (anonymous):

I did but my could not get the same answer that my prof. gave me. His answer (1,-1,-2) I tested his answer but the last equation seems not work out

OpenStudy (anonymous):

can you show me the solution of your professor?

OpenStudy (anonymous):

By the way, I missing, let me retype my question: 2A+3B-4C=7 A-5B+2C=2 -7A-3B+5C=-20 and his answer is (1,-1,-2)

OpenStudy (anonymous):

try this: equation 1 minus equation 2, equation 1 minus equation 3 and equation 2 minus equation 3 1-2: by elimination 13B-8C=3 1-3: -5A+C=27 2-3: 38B+19C=-6